OpenStudy (anonymous):

A hydrate of CoCl2 has a mass of 15.80 g before heating. After heating, the mass of the anhydrous compound is found to be 9.32 g. Explain how you would determine the formula of the hydrate and then write out the full name of the hydrate. Please help?!?!

OpenStudy (anonymous):

I dont know how to do this...and Im in the middle of a test. I dont want to cheat-I need help on how to do this.

OpenStudy (anonymous):

How would I set up this equation?

OpenStudy (callisto):

Let the formula be \[CoCl_{2}. nH_{2}O\]\[CoCl_{2}. nH_{2}O->CoCl_{2} .nH_{2}O\] Mass of H2O = 15.8-9.32 = 6.48g no. of mole of H2O = 6.48/(1.0x2+16.0) = 0.36 No. of mole of CoCl2 = 9.32/(58.9+35.5x2) = 0.071747 n = 0.36/0.071747 = 5 I think it should be like that. Hope it helps

OpenStudy (callisto):

-> CoCl2 + nH2O , sorry typing mistake

OpenStudy (anonymous):

That did help, i see it now..thanks so much:)

OpenStudy (callisto):

When a hydrated salt is heated, the water would loss i mean not attached to the compound Then, let the formula of the hydrated compound be ____ . n H2O As the water is 'detached' after heating, just separate the 2, and you'll come up with the formula, that is _____ . nH2O -> _____ + nH2O then, calculate the number of moles and find their ratio. => done! :)

OpenStudy (anonymous):

Thank you, thank you, thank you! I was stuck on this question for like 10 mins. before I came here..lol

OpenStudy (callisto):

Welcome! :)